OpenStudy (anonymous):

The function for the cost of materials to make a muffin is f(x) = Three fourths x + 3, where x is the number of muffins. The function for the selling price of those muffins is g(f(x)), where g(x) = 3x + 4. Find the selling price of 20 muffins.

OpenStudy (gorv):

so x=20 we have to find selling price=g(f(x)) for 20 muffins selling price=g(f(20) f(20)=3/4 *20+3 =3*5+3=18

OpenStudy (gorv):

selling price=g(f(x))=g(f(20))=g(18)=3*20+4=60+4 =64